# Pointing the Way: Refining Radar-Lidar Localization Using Learned ICP Weights

### v2

@@ Abstract
This paper presents a novel deep-learning-based approach to improve localizing radar measurements against lidar maps. Although the state of the art for localization is matching lidar data to lidar maps, radar has been considered as a promising alternative. This is largely due to radar being more resilient against adverse weather such as precipitation and heavy fog. To make use of existing high-quality lidar maps, while maintaining performance in adverse weather, it is of interest to match radar data to lidar maps. However, owing in part to the unique artefacts present in radar measurements, radar-lidar localization has struggled to achieve comparable performance to lidar-lidar systems, preventing it from being viable for autonomous driving. This work builds on an ICP-based radar-lidar localization system by including a learned preprocessing step that weights radar points based on high-level scan information. Combining a proven analytical approach with a learned weight reduces localization errors in radar-lidar ICP results run on real-world autonomous driving data by up to $54.94\%$ in translation and $68.39\%$ in rotation, while maintaining interpretability and robustness.
@@ I INTRODUCTION
The current state of the art for the task of localization is matching live lidar scans to preconstructed lidar maps [1]. Lidar sensors are capable of capturing high-definition maps, which can then be used to achieve accurate localization. However, lidar measurements can be affected or completely blocked by precipitation and other challenging environmental conditions [2, 3, 4, 5]. Radar has been considered as an alternative for localization in all-weather conditions. Radar, owing to its longer wavelength, is much less susceptible to small particles in the air, and is thus largely unaffected by rain, snow, dust, or fog [6]. However, the longer wavelengths and lower resolution inherently lead to sparser maps as compared to those constructed using lidar. Additionally, radar measurements suffer from artefacts such as speckle (noisy measurements), saturation (falsely strong return signal), and ghosting (multiple echoes of the same point or object) [7]. These artefacts are challenging to model and heuristically extract [8]. As a result, radar-radar localization accuracy is still worse than that of lidar-lidar [2], making it not quite viable for tasks that require centimeter-level accuracy, such as autonomous driving.
Figure 1: A satellite image of the environment, overlaid with the full extracted radar pointcloud (left) and the radar pointcloud weighted using our method (right). The weight associated with each point in the weighted pointcloud corresponds to the intensity in colour. It can be seen that our method picks out some of the most important geometric features while ignoring noisy and unreliable points, thus facilitating the task of ICP.
Owing to the unique advantages and disadvantages of lidar and radar, there has been interest in multi-modal radar-lidar localization. By localizing using radar scans against already available high-quality lidar maps, the goal is to achieve near state-of-the-art localization performance in all weather conditions. However, past work in radar-lidar localization has not been able to achieve the accuracy required for autonomous driving [2, 9, 5, 10]. Our previous work, [2], was able to achieve state-of-the-art radar-lidar localization by using the iterative closest point (ICP) algorithm to match radar and lidar pointclouds, but still fell slightly short of desired localization accuracy [11].
@@ II RELATED WORK
Radar has seen a resurgence within robotics in recent years [12, 13, 14]. The ability of modern-day radar systems to produce reasonable long-term odometry results is shown in [15], where an ICP-style pointcloud alignment is used to compute the relative motion between consecutive timestamps. An alternative approach is presented in [16], where a network is trained to mask raw radar scans before using them for correlative scan matching of consecutive frames. It is shown that such a learned approach is capable of outperforming heuristic filtering methods. A network that directly extracts features, descriptors, and feature scores from raw radar images and then uses them to match radar data from consecutive timesteps is presented in [17]. In our previous work [18], features are learned in a self-supervised manner, without needing groundtruth data. Although feature-extraction methods are typically more interpretable and robust to poor initial guesses, they frequently require large networks with higher computational cost. Additionally, feature matching across sensing modalities, such as between radar and lidar, poses an extra challenge. Finally, it is important to note that most radar algorithms have focused on odometry or simultaneous localization and mapping (SLAM), e.g. [19, 20, 21, 22], meaning that investigations into the performance of radar in localization against a previously known map is understudied. An extended Kalman filter (EKF) ICP-based approach for radar-radar localization is presented in [23]. However, their localization errors are on the order of meters, far above autonomous driving requirements. In our recent ICP-based approach [2], localization is done using batch optimization with a modified Constant False Alarm Rate (CFAR) [24] point extractor called Bounded False Alarm Rate (BFAR) [25]. This work shows that radar localization against an existing radar map can achieve promising results that approach the level needed for autonomous driving.
@@
Radar has also been used to localize against lidar maps. One of the first algorithms for radar-lidar localization was proposed for indoor disaster environments where lidar is blocked by smoke [5]. This algorithm implements a recursive undistortion approach to utilize shape matching between live radar scans and an existing lidar map. This recursive implementation is shown to overcome some aspects of radar noise and artefacts to achieve usable localization results in situations where lidar measurements completely fail. A learning-based approach was taken in [26], where a generative adversarial network (GAN) was used to style-transfer a live radar scan to lidar data. Instead of doing a cross-sensor style transfer, Yin et al. [9] use a set of networks to transfer both radar and lidar scans to a common learned embedding, in which they then perform cross-correlation matching. This match is then treated as a measurement in a Kalman-filter framework. Their algorithm is set up to be fully differentiable, meaning that the networks can be trained to produce an embedding that directly minimizes the final localization output from the Kalman filter. Another embedding-style approach is presented in [10], where it is hypothesized that although extracted radar and lidar points do not necessarily have one-to-one correspondences, their groupings are related. Thus, radar points are binned into discrete Cartesian and polar groupings and an initial alignment between radar and lidar groupings is done via a density metric. The estimate is then further refined using ICP. A learned radar-lidar place recognition algorithm is presented in [27], but they do not evaluate localization accuracy. Although promising, none of these papers are able to achieve centimeter-level localization accuracy, a requirement for safe autonomous vehicle operations. However, our previous work was able to achieve an accuracy on the order of tens of centimeters with a nonlearned approach [2]. This algorithm first converts the radar and lidar measurements to pointclouds using BFAR, and then applies ICP to match the pointclouds.
@@
In our previous work, we have found that CFAR-based point detectors work well for generating radar pointclouds. These can subsequently be matched to prebuilt lidar maps using ICP to yield state-of-the-art radar-lidar performance. This approach makes use of the full-resolution radar scan without needing to downsample, as in [18], and does not require any discretization of the pose estimate, as in [9, 16]. However, CFAR-like detectors only use local information to extract points, meaning that a significant amount of noise gets extracted. Thus, our approach is to start with the radar-lidar localization pipeline presented in [2] and, building on [16], incorporate a learned weight mask into the ICP algorithm to identify only the most useful radar points for the task of radar-lidar localization. Using a network allows us to learn the high-level spatial contextual cues in order to softly reject points that may correspond to noise, while still retaining the high-resolution, discretization-free point localization from CFAR and ICP.
@@ III METHODOLOGY
This section presents the various parts of the pipeline. First, the overall training architecture is presented. Then, the differentiable ICP (dICP) and supervisory map mask components are discussed in more detail. Our code base is publicly available to facilitate reproducibility and extensions [^1] Paper code is available at: https://github.com/lisusdaniil/mm_masking..
@@
Figure 2: The pipeline used to train the weight mask network. A radar scan, used to localize against a reference lidar map, is fed to a U-Net network that produces a weight mask. In parallel, the scan is used to extract a pointcloud using the BFAR detector. The pointcloud then indexes weights from the weight mask. The points and corresponding weights are used in a weighted differentiable ICP algorithm (dICP) to localize against the reference map pointcloud. The training loss is formed through a term corresponding to the error in the dICP result and a binary cross-entropy (BCE) term computed between the weight mask and a supervisory map mask. The network is trained by backpropagating the loss through dICP and the weight extraction.
III-A Architecture
@@
The overall architecture is presented in Figure 2. The input to the pipeline is a single radar intensity scan. A Cartesian representation of the scan, normalized to have a maximum pixel value of one, is used as input to the network, whereas an unnormalized polar representation is fed to the point-extraction module. Inspired by [16], we use a U-Net network [28] to learn the weight mask for each radar input. A U-Net lends itself well to our application, as we are, in essence, trying to segment the radar scan into ‘good’ and ‘bad’ areas for pointcloud extraction. The encoder increases the number of channels from one, the measured intensity, to 8, 16, 32, 64, 128, and 256 channels. Each channel increase is composed of convolution blocks that include a 2D convolutional layer (kernel size 3, padding 1), a ReLU activation, another 2D convolution, a dropout layer with probability of 0.05, and finally a max-pooling operation (kernel size 2, stride 2). Using dropout helps to prevent over-fitting and encourages the network to learn higher-level, general information instead of relying on specific pixels [29]. The decoder then decreases the channel number from 256 down to 8 using the same steps as the encoder. For each decrease, the input is up-sampled, passed through the same convolution block as described for the encoder, concatenated with the corresponding skip connection from the encoder section, and passed through a convolution block again. Finally, the output is passed through one last 2D convolutional layer to decrease the 8 channels back down to one, and then through a sigmoid activation. The resulting mask is then normalized to have a maximum value of one. The network is implemented in PyTorch [30].
@@
Note, that as radar is a 2D sensor, the overall pipeline runs in 2D. This means that the dICP algorithm estimates a transform from the radar scan to the lidar map, ${\hat{\mbf{T}}}_{lr}\in SE(2)$, which is contrasted with the groundtruth pose, $\mbf{T}_{lr}\in SE(2)$, to form the error vector $\mbf{e}$ according to
$\mbf{e}=\begin{bmatrix}{e}_{x}&{e}_{y}&{e}_{\phi}\end{bmatrix}^{\mathsf{T}}=\log\left({\hat{\mbf{T}}}_{lr}\mbf{T}_{lr}^{-1}\right)^{\vee},$ (2)
where $\log(\cdot)$ is the logarithmic operator that converts $SE(2)$ elements to their Lie algebra representation and $(\cdot)^{\vee}$ is the operator that converts the Lie algebra representation to $\mathbb{R}^{3}$ [31]. Additionally, ${e}_{x}$, ${e}_{y}$, and ${e}_{\phi}$ are the $x$ (longitudinal), $y$ (lateral), and $\phi$ (heading) error components, respectively.
The loss is then formed based on two signals. The first, $\mathcal{L}_{\mathrm{ICP}}$, is based on $\mbf{e}$, and is formed as
$\mathcal{L}_{\mathrm{ICP}}=\mbf{e}^{\mathsf{T}}\begin{bmatrix}\alpha&0&0\\ 0&\alpha&0\\ 0&0&\beta\end{bmatrix}\mbf{e},$ (6)
where $\alpha,\beta\in\mathbb{R}$ control the relative impact of each pose term.
The second loss signal, $\mathcal{L}_{\mathrm{BCE}}$, is a binary cross-entropy loss computed between the generated weight mask and a supervisory map mask discussed in Section III-C. The final loss term is then of the form
@@
where $\gamma\in\mathbb{R}$ controls the relative impact of $\mathcal{L}_{\mathrm{BCE}}$ on $\mathcal{L}$.

### v3

@@ Abstract
This paper presents a novel deep-learning-based approach to improve localizing radar measurements against lidar maps. This radar-lidar localization leverages the benefits of both sensors; radar is resilient against adverse weather, while lidar produces high-quality maps in clear conditions. However, owing in part to the unique artefacts present in radar measurements, radar-lidar localization has struggled to achieve comparable performance to lidar-lidar systems, preventing it from being viable for autonomous driving. This work builds on ICP-based radar-lidar localization by including a learned preprocessing step that weights radar points based on high-level scan information. To train the weight-generating network, we present a novel, stand-alone, open-source differentiable ICP library. The learned weights facilitate ICP by filtering out harmful radar points related to artefacts, noise, and even vehicles on the road. Combining an analytical approach with a learned weight reduces overall localization errors and improves convergence in radar-lidar ICP results run on real-world autonomous driving data. Our code base is publicly available to facilitate reproducibility and extensions222 Paper code is available at: https://github.com/utiasASRL/mm_masking..
@@ I INTRODUCTION
The current state of the art for the task of localization is matching live lidar scans to preconstructed lidar maps [survey_lidar_loc]. Lidar sensors are capable of capturing high-definition maps, which can then be used to achieve accurate localization. However, lidar measurements can be affected or completely blocked by precipitation and other challenging environmental conditions [are_we_ready_for, yoneda2019automated, Courcelle_Baril_Pomerleau_Laconte_2022, Park_Kim_Kim_2019]. Radar has been considered as an alternative for localization in all-weather conditions. Radar, owing to its longer wavelength, is much less susceptible to small particles in the air, and is thus largely unaffected by rain, snow, dust, or fog [weather_sensor_impact]. However, the longer wavelengths and lower resolution inherently lead to sparser maps as compared to those constructed using lidar. Additionally, radar measurements suffer from artefacts such as speckle (noisy measurements), saturation (falsely strong return signal), and ghosting (multiple echoes of the same point or object) [radar_nav_book]. These artefacts are challenging to model [Weston_Jones_Posner_2020]. As a result, radar-radar localization accuracy is still worse than that of lidar-lidar [are_we_ready_for], making it not viable for tasks that require centimeter-level accuracy, such as autonomous driving.
Refer to captionWeighted Radar ScanFull Radar ScanLidar MapEcho removedNoise removedWalls highlighted
Figure 1: A satellite image of the environment, overlaid with the radar pointcloud weighted using our method (red, top), full extracted radar pointcloud (red, bottom), and lidar map (blue). The weight associated with each point in the weighted pointcloud corresponds to the intensity in colour. Our method picks out most of the important geometric features while ignoring noisy points and radar artefacts, thus facilitating the task of ICP.
Owing to the unique advantages and disadvantages of lidar and radar, there has been interest in multi-modal radar-lidar localization. By localizing using radar scans against already available high-quality lidar maps, the goal is to achieve near state-of-the-art localization performance, while being resilient to all weather conditions. However, past work in radar-lidar localization has not been able to achieve the accuracy required for autonomous driving [are_we_ready_for, RaLL, Park_Kim_Kim_2019, RoLM]. Our previous work, [are_we_ready_for], was able to achieve state-of-the-art radar-lidar localization by using the iterative closest point (ICP) algorithm to match radar and lidar pointclouds, but still fell short of desired localization accuracy [icp_ref].
@@ II RELATED WORK
Radar has seen a resurgence within robotics in recent years [radar_survey, radar_survey_2, venon2022millimeter]. The ability of modern-day radar systems to produce reasonable long-term odometry results is shown in [Cen_Newman_2018], where an ICP-style pointcloud alignment is used to compute the relative motion between consecutive timestamps. An alternative approach is presented in [masking_by_moving], where a network is trained to mask raw radar scans before using them for correlative scan matching of consecutive frames. It is shown that such a learned approach is capable of outperforming heuristic filtering methods. A network that directly extracts features, descriptors, and feature scores from raw radar images and then uses them to match radar data from consecutive timesteps is presented in [under_the_radar]. A self-supervised feature learning approach, which does not require groundtruth data, is shown in [hero_paper]. Although feature-extraction methods are typically more interpretable and robust to poor initial guesses, they frequently require large networks with higher computational cost. Additionally, feature matching across sensing modalities, such as between radar and lidar, poses an extra challenge. Finally, most radar algorithms have focused on odometry or simultaneous localization and mapping (SLAM), e.g., [Jose_Adams_2004, Rouveure_Monod_Faure_2009, Checchin_Gerossier_Blanc_Chapuis_Trassoudaine_2010, Schuster_Keller_Rapp_Haueis_Curio_2016], meaning that investigations into the performance of radar in localization against a previously known map is understudied. An extended Kalman filter (EKF) ICP-based approach for radar-radar localization is presented in [Ward_Folkesson_2016]. However, their localization errors are on the order of meters, far above autonomous driving requirements. In our recent ICP-based approach [are_we_ready_for], localization is performed using batch optimization with a modified Constant False Alarm Rate (CFAR) [cfar] point extractor called Bounded False Alarm Rate (BFAR) [bfar]. This work shows that radar localization against an existing radar map can approach the level needed for autonomous driving, with translation errors of $6-10\,$\mathrm{cm}$$.
@@
Radar has also been used to localize against lidar maps. One of the first algorithms for radar-lidar localization was proposed for indoor disaster environments where lidar is blocked by smoke [Park_Kim_Kim_2019]. This algorithm implements a recursive undistortion approach to utilize shape matching between live radar scans and an existing lidar map. This recursive implementation is shown to overcome some aspects of radar noise and artefacts to achieve usable localization results in situations where lidar measurements completely fail. A learning-based approach was taken in [radar_on_lidar], where a generative adversarial network (GAN) was used to style-transfer a live radar scan to lidar data. Instead of doing a cross-sensor style transfer, Yin et al. [RaLL] use a set of networks to transfer both radar and lidar scans to a common learned embedding in which they then perform cross-correlation matching. This match is then treated as a measurement in a Kalman-filter framework. Their algorithm is set up to be fully differentiable, meaning that the networks can be trained to produce an embedding that directly minimizes the final localization output from the Kalman filter. Another embedding-style approach is presented in [RoLM], where it is hypothesized that although extracted radar and lidar points do not necessarily have one-to-one correspondences, their groupings are related. Thus, radar points are binned into discrete Cartesian and polar groupings and an initial alignment between radar and lidar groupings is done via a density metric. The estimate is then further refined using ICP. A learned radar-lidar place recognition algorithm is presented in [Yin_Xu_Wang_Xiong_2021], but they do not evaluate localization accuracy. Although promising, none of these papers are able to achieve centimeter-level localization accuracy, a requirement for safe autonomous vehicle operations. However, our previous work was able to achieve an accuracy on the order of tens of centimeters with a nonlearned approach [are_we_ready_for]. This algorithm first converts the radar and lidar measurements to pointclouds using BFAR, and then applies ICP to match the pointclouds.
@@
In our previous work, we have found that CFAR-based point detectors work well for generating radar pointclouds. These can subsequently be matched to prebuilt lidar maps using ICP to yield state-of-the-art radar-lidar performance. This approach makes use of the full-resolution radar scan without needing to downsample, as in [hero_paper], and does not require any discretization of the pose estimate, as in [RaLL, masking_by_moving]. However, CFAR-like detectors only use local information to extract points, meaning that a significant amount of noise gets extracted. Thus, our approach is to start with the radar-lidar localization pipeline presented in [are_we_ready_for] and, building on [masking_by_moving], incorporate a learned weight mask into the ICP algorithm to identify only the most useful radar points for the task of radar-lidar localization. Using a network allows us to learn the high-level spatial contextual cues in order to softly reject points that may correspond to noise, while still retaining the high-resolution, discretization-free point localization from CFAR and ICP.
@@ III METHODOLOGY
Figure 2: The pipeline used to train the weight mask network. A radar scan is fed to a U-Net network that produces a weight mask. In parallel, the scan is used to extract a pointcloud using the BFAR detector. The pointcloud then indexes weights from the weight mask. The points and corresponding weights are used in a weighted differentiable ICP algorithm (dICP) to localize against the reference map pointcloud. The training loss is formed from the error in the dICP result and a binary cross-entropy (BCE) term computed between the weight mask and a supervisory map mask. The network is trained by backpropagating the loss through dICP and the weight extraction.
III-A Weighted Point-to-point ICP
This section presents the specific ICP implementation that our pipeline aims to improve. The goal of ICP is to align an $n$-point, $D$-dimensional source pointcloud $\mbf{S}\in\mathbb{R}^{n\times D}$ to an $m$-point, $D$-dimensional target pointcloud $\mbf{P}\in\mathbb{R}^{m\times D}$. ICP solves this alignment problem in an iterative fashion to find the transformation $\mbf{T}_{ts}\in SE(D)$, where $SE(D)$ is the special Euclidean group of dimension $D$ [barfoot2017state]. In our case, the source pointcloud is the radar scan, the target pointcloud is the lidar map, and $D=2$.
First, for each point $\mbf{s}_{i}$ in $\mbf{S}$, we find the nearest neighbour $\mbf{q}_{i}$ from $\mbf{P}$. The definition of ‘nearest’ is a design choice, and we only present the so-called point-to-point definition as it is the one used in this paper. However, our method could be easily be applied to other common definitions, for example point-to-plane, without substantial changes. In the case of point-to-point, we compute $\mbf{q}_{i}$ as
$\mbf{q}_{i}=\mathrm{argmin}_{\mbf{q}\in\mbf{P}}\left(\left\|{\check{\mbf{T}}}_{ts}\mbf{s}_{i}-\mbf{q}\right\|_{2}\right)\in\mathbb{R}^{D},$ (1)
where we use the latest estimate ${\check{\mbf{T}}}_{ts}$ and abuse notation to have ${\check{\mbf{T}}}_{ts}\mbf{s}_{i}\in\mathbb{R}^{D}$. An error is then formed between $\mbf{s}_{i}$ and $\mbf{q}_{i}$ as
$\mbf{e}_{i}={\check{\mbf{T}}}_{ts}\mbf{s}_{i}-\mbf{q}_{i}\in\mathbb{R}^{D}.$ (2)
A robust-cost function $\rho(\cdot)\colon\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ can be used to down-weight large outlier errors by re-scaling the error as $\rho(\mbf{e}_{i})$ [robust_func]. A trimmed ICP modification, which we use here, can be included into $\rho(\cdot)$ to zero-out error terms that have a magnitude above some threshold. If prior knowledge about the reliability of specific points as compared to others exists, a weight $\mbf{w}_{p}\in\mathbb{R}^{n\cdot D}$ is included in the optimization problem. The final objective function is then
$J(\mbf{T}_{ts})=\frac{1}{2}\mbf{e}^{\mathsf{T}}\mbf{W}\mbf{e}\in\mathbb{R},$ (3)
where $\mbf{e}\in\mathbb{R}^{n\cdot D}$ is a stacked vector of all $\mbf{e}_{i}$, $\rho(\mbf{e})$ is applied error-wise for an element-wise multiplication $\odot$ in $\mbf{W}=\mathrm{diag}(\rho(\mbf{e})\odot\mbf{w}_{p})\in\mathbb{R}^{n\cdot D\times n\cdot D}$, and which can be minimized using least-squares to find an optimal ${\hat{\mbf{T}}}_{ts}$. This paper aims to find an optimal $\mbf{w}_{p}$.
III-B Architecture
@@
Our overall pipeline architecture is presented in Figure 2. The input to the pipeline is a single radar intensity scan. A Cartesian representation of the scan, normalized to have a maximum pixel value of one, is used as input to the network, whereas an unnormalized polar representation is fed to the point-extraction module. Inspired by [masking_by_moving], we use a U-Net network [unet] to learn the weight mask for each radar input. A U-Net lends itself well to our application, as we are, in essence, trying to segment the radar scan into ‘good’ and ‘bad’ areas for pointcloud extraction. The encoder increases the number of channels from one, the measured intensity, to 8, 16, 32, 64, 128, and 256. Each channel increase is composed of convolution blocks that include a 2D convolutional layer (kernel size 3, padding 1), a ReLU activation, another 2D convolution, a dropout layer with probability of 0.05, and finally a max-pooling operation (kernel size 2, stride 2). Using dropout helps to prevent over-fitting and encourages the network to learn higher-level, general information instead of relying on specific pixels [dropout]. The decoder then decreases the channel number from 256 down to 8 using the same steps as the encoder. For each decrease, the input is up-sampled, passed through the same convolution block as described for the encoder, concatenated with the corresponding skip connection from the encoder section, and passed through a convolution block again. Finally, the output is passed through one last 2D convolutional layer to decrease the 8 channels back down to one, and then through a sigmoid activation. The resulting mask is then normalized to have a maximum value of one. The network is implemented in PyTorch [pytorch].
@@
Note, that as radar is a 2D sensor, the overall pipeline runs in 2D. This means that the dICP algorithm estimates a transform from the radar scan to the lidar map, ${\hat{\mbf{T}}}_{lr}\in SE(2)$, which is contrasted with the groundtruth pose, $\mbf{T}_{lr}\in SE(2)$, to form the ICP error vector $\mbf{e}_{\mathrm{ICP}}$ according to
$\mbf{e}_{\mathrm{ICP}}=\begin{bmatrix}{e}_{x}&{e}_{y}&{e}_{\phi}\end{bmatrix}^{\mathsf{T}}=\log\left({\hat{\mbf{T}}}_{lr}\mbf{T}_{lr}^{-1}\right)^{\vee},$ (4)
where $\log(\cdot)$ is the logarithmic operator that converts $SE(2)$ elements to their Lie algebra representation and $(\cdot)^{\vee}$ is the operator that converts the Lie algebra representation to $\mathbb{R}^{3}$ [barfoot2017state]. Additionally, ${e}_{x}$, ${e}_{y}$, and ${e}_{\phi}$ are the $x$ (longitudinal), $y$ (lateral), and $\phi$ (heading) error components, respectively.
The loss is then formed based on two signals. The first, $\mathcal{L}_{\mathrm{ICP}}$, is based on $\mbf{e}_{\mathrm{ICP}}$, and is formed as
$\mathcal{L}_{\mathrm{ICP}}=\mbf{e}_{\mathrm{ICP}}^{\mathsf{T}}\begin{bmatrix}\alpha&0&0\\ 0&\alpha&0\\ 0&0&\beta\end{bmatrix}\mbf{e}_{\mathrm{ICP}},$ (5)
where $\alpha,\beta\in\mathbb{R}_{\geq 0}$ control the relative impact of each translation and rotation term.
The second loss signal, $\mathcal{L}_{\mathrm{BCE}}$, is a binary cross-entropy loss computed between the generated weight mask and a supervisory map mask discussed in Section III-D. The final loss term is then of the form
@@
where $\gamma\in\mathbb{R}_{\geq 0}$ controls the relative impact of $\mathcal{L}_{\mathrm{BCE}}$ on $\mathcal{L}$.